which of the following is greater than 16/19
a.15/18
b.17/20
c.14/17
d.13/16
1
Expert's answer
2013-09-12T07:42:31-0400
Calculate the difference, reducing fractions to a common denominator.
a. 15/18 - 16/19 = 285/342 - 288/342 = -3/342 <0
b. 17/20 - 16/19 = 323/380 - 320/380 = 3/380 >0
c. 14/17 - 16/19 = 266/323 - 272/323 = -6/323 <0
d. 13/16 - 16/19 = 247/304 - 256/304 = -9/304 <0
Answer: b. 17/20